### v4

# Yang-Mills Field in the $\kappa$-space-time

Abstract
In this paper, we construct Yang-Mills theory in the $\kappa$-space-time. We derive Yang-Mills equations in the $\kappa$-deformed space-time, valid up to first order in the deformation parameter $a$, using the generalisation of Feynman’s approach. We first derive the $\kappa$-deformed Wong’s equation and using it, in the Jacobi identity involving velocities and coordinates of $\kappa$-deformed space-time, we derive the $\kappa$-deformed homogeneous Yang-Mills equations. We then construct the Lagrangian for Yang-Mills theory in $\kappa$-deformed space-time and show that it is invariant under SU(N) transformation. We also derive the expression for the force experienced by an isospin-carrying particle in the presence of Yang-Mills field in the $\kappa$-space-time.
1 Introduction
Microscopic theory of gravity requires amalgamation of general relativity and quantum mechanics and such a theory is needed to understand and explain the nature of gravitational force at very short distances, well below nuclear size. Though many different approaches have been developed to discuss quantum gravity, with varying degrees of success, a completely satisfactory quantum theory of gravity still remains elusive. But one of the common features all these approaches share is the existence of a fundamental length scale; it is below this length scale, the quantum effects of gravity become important. As non-commutative space-time(s) inherently introduce a length scale, non-commutative geometry promises to be a viable framework for consistent quantisation of gravity [1]. Thus it is also interesting to construct and study field theory models on non-commutative (NC) space-times and analyse their quantization. In particular, construction of standard model in NC space-time is of immense importance and this was carried out in [1, 2]. These studies naturally led to the construction and investigation of gauge theories on NC space-time. NC standard model along with a consistent quantum gravity also opens up the possibility of grand unified models [1, 2].
As the existence of a fundamental length scale is not compatible with the special theory of relativity, principle of relativity needed to be modified. This lead to doubly special relativity (DSR) which admits a fundamental length scale and speed of light as universal constants [3]. The underlying space-time of DSR is known to be $\kappa$-space- time [4, 5, 6]. This $\kappa$-space-time also appears in the low energy limit of quantum gravity models [7]. The symmetries of $\kappa$-space-time is realised as Hopf algebra [4, 5]. Various issues in constructing quantum field theories in $\kappa$-space-time are addressed and resolved in [4, 5, 6, 7, 8, 9, 10, 11, 12, 16, 17, 18, 13, 14, 15, 19, 20, 21]
The present status of research on building gauge theories in various non-commutative space-times as well as attempts to include gravity into NC framework is discussed in [22]. Deformed gauge theories are constructed for arbitrary compact Lie groups and the Lagrangian for such gauge theories possessing a deformed Lorentz covariance are constructed in [13]. The symmetries of $\kappa$-space-time is realised using undeformed $\kappa$-Poincare algebra [23]. Here different possible realisations of space-time coordinates of $\kappa$-space-time in terms of the commutative variables are introduced.
The studies of quantum field theories in $\kappa$-space time is carried out in [27, 24, 25, 26, 28, 29]. Twisted flip operator which is relevant for describing particle statistics in $\kappa$-space-time is obtained in [24]. In [25], deformed statistics and oscillator algebra of quantum fields in $\kappa$-space-time are studied. In [26], the Dirac equation in $\kappa$-Minkowski space-time is constructed and it was shown that the charge conjugation invariance is lost while preserving parity and time reversal symmetries. In these studies, $\kappa$-deformed space-time coordinates were written in terms of commutative coordinates and their derivatives. This provided a systematic way of writing down field theory models in $\kappa$-space-time in terms of commutative variables. This approach has the advantage of applying well established field theoretic tools of commutative space-time in analysing the effects of $\kappa$-deformation [23].
In [30], Maxwell’s equations were derived by assuming Newton’s law and commutation relations between position coordinates and corresponding velocities. Repeated use of Jacobi identity involving coordinates, velocities, and different combinations of coordinates and velocities then lead to two of the Maxwell’s equations. The remaining equations were argued to be defining the sources of electric and magnetic fields, respectively, [30]. This was followed by the derivation of Yang-Mills equation by considering a non-relativistic particle carrying isospin [31]. In[32], a covariant generalisation of the derivation of the Maxwell’s equations and Yang-Mills equation were developed. The construction of Yang-Mills equation in this approach crucially depends on the Wong’s equation which describes the motion of a particle carrying isospin charge, in the presence of external non-abelian gauge field [33].
Maxwell’s equation in $\kappa$-space-time was derived by generalising Feynman’s approach in [27]. Using the minimal coupling method, $\kappa$-deformed Maxwell’s equation was derived in [28]. In this approach, corrections to Maxwell’s equations, valid up to first order in the deformation parameter were derived. Here the deformed Maxwell’s equations were written in terms of fields defined on the commutative space-time. Generalising this minimal coupling prescription to curved, $\kappa$-deformed space-time, geodesic equation in $\kappa$-deformed space-time was derived and analysed in [29].
Construction and analysis of non-abelian gauge theories in $\kappa$-space-time is an active area of research [13, 9, 10, 11, 12, 18, 16, 15, 14, 17]. In [14, 15], a general formalism for constructing field theories on quantum space-time is developed and used to obtain field theory models on $\kappa$-space-time. This was followed by the construction of U(1) gauge theory in $\kappa$-space-time using twist formalism [16]. U(1) gauge theory on $\kappa$-Minkowski space is discussed in[17]. The complications brought by the interconnection between translations and gauge symmetries in NC space-times in constructing Hodge-dual-field-strength tensor were addressed in [18]. Using Seiberg-Witten map, the gauge theory on kappa-space-time was related to a theory defined on the commutative space-time having corrections depending on the deformation parameter. NC gauge theory has been constructed in $\kappa$-Minkowski space-time by applying twist formalism in [19].
In this work, we generalise the Feynman’s approach in its covariant form to derive $\kappa$-deformed Yang-Mills equation. We also obtain the corresponding $\kappa$-deformed Yang-Mills Lagrangian. The $\kappa$-deformed Yang-Mills equation is derived by exploiting Jacobi identities involving various combinations of $\kappa$-deformed coordinates and corresponding velocities. In this approach, taking into account of the anti-symmetry property of the commutators between velocities, we introduce a second rank tensor, which is later shown to be the Yang-Mills field strength. The Jacobi identities then show that this Yang-Mills field strength satisfies homogeneous Yang-Mills equations. In evaluating the Jacobi identities, one needs the commutation relations between $\kappa$-deformed velocities with Lie algebra valued functions defined on $\kappa$-space-time. This naturally requires the commutation relation between different components of the $\kappa$-deformed velocities with generators of the Lie algebra and their time derivatives. The Wong’s equation, describing the motion of isospin carrying particle in the background of non-abelian gauge field is used in the derivation of Yang-Mills equation at this stage. The Wong’s equation comes with a condition on the time development of the generators of the Lie algebra associated with the non-abelian gauge field. This condition plays an important role in the derivation of Yang-Mills field equation using the Feynman’s approach.
Here we derive Yang-Mills equation in $\kappa$-space-time whose coordinate satisfy
$[\hat{x}_{\mu},\hat{x}_{\nu}]=i(a_{\mu}\hat{x}_{\nu}-a_{\nu}\hat{x}_{\mu}).$ (1.1)
with $a_{0}=a=\frac{1}{\kappa}$, $a_{i}=0$, by enlarging the scope of covariant generalisation of the Feynman’s approach to NC space-time. This requires the derivation of $\kappa$-deformed Wong’s equation. Using the $\kappa$-deformed Dirac Hamiltonian describing the interaction of isospin carrying fermion with non-abelian gauge field, we obtain the $\kappa$-deformed Wong’s equation. With the aid of this, we find the commutation relation between $\kappa$-deformed velocity and a Lie algebra valued functions in the $\kappa$-space-time. This also serves to provide the explicit form of gauge covariant derivative. Now replacing $[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]$ with a second rank, anti-symmetric tensor, in the Jacobi identity involving $\kappa$-deformed velocities, we obtain the equation satisfied by components of this second rank tensor. Employing Jacobi identities involving the deformed space-time coordinates and the generator of the Lie algebra, we find the second rank tensor introduced to be the Yang-Mills field strength in the $\kappa$-deformed space-time. Thus we derive the “homogeneous” Yang-Mills equation in the $\kappa$-deformed space-time. Using the $\kappa$-deformed field strength, we then write down the Lagrangian describing Yang-Mills field theory in the $\kappa$-deformed space-time. The remaining Yang-Mills equations follow as Euler-Lagrange equations from this $\kappa$-deformed Lagrangian. The force equation for a particle carrying isospin, interacting with Yang-Mills field is then derived.
The derivation of $\kappa$-deformed Wong’s equation also provides definition of gauge covariant derivative in $\kappa$-deformed space-time. We show that the commutators of this $\kappa$-deformed gauge covariant derivatives exactly match with the deformed Yang-Mills field strength we derived in the $\kappa$-space-time. We see that the gauge covariant derivatives and hence non-abelian “electric” and “magnetic” fields in the $\kappa$-space time are modified in the same general form but different in the numerical factors.
After quickly recalling the derivations of Maxwell’s equations and Yang-Mills equation using Feynman’s approach, we present the derivation of $\kappa$-deformed Wong’s equation. We then derive Yang-Mills equation in the $\kappa$-deformed space-time. Commutator between the $\kappa$-deformed velocity and Lie algebra valued functions of $\kappa$-deformed space-time is an integral part of this derivation and is calculated using $\kappa$-deformed Wong’s equation. Using the explicit form of the Yang-Mills field strength in the $\kappa$-deformed space-time obtained, we then construct the Lagrangian describing $\kappa$-deformed Yang-Mills theory. We show that this model is invariant under SU(N) transformations. We then derive the force experienced by an isospin carrying particle moving in presence of Yang-Mills fields.
2 Gauge field equations from Newtons’s Law: Feynman’s approach
In Feynman’s approach [30], Maxwell’s equations are derived by assuming Newton’s law and the commutation relations between position coordinates and velocities, i.e.,
$m\ddot{x}_{j}=F_{j}(x,\dot{x},t)$ (2.1)
$[x_{i},x_{j}]=0~{}~{}~{}m[x_{j},\dot{x}_{k}]=i\hbar\delta_{jk}.$ (2.2)
Using time derivative of the commutator given above and Jacobi identity between $x_{i}$ and velocities, one finds $[x_{i},F_{j}]$ to be antisymmetric in indices. Thus an anti-symmetric tensor is introduced as
$[x_{j},F_{k}]=-\frac{i\hbar}{m}\epsilon_{jkl}B_{l}$ (2.3)
Considering eq.(2.3) as the definition of B and using Jacobi identity, it is shown that B is function of x and t only. Use of eq.(2.2) in eq.(2.3) allows one to find $F_{k}$ by integrating right hand side of eq.(2.3) with respect to $\dot{x}_{k}$ setting
$F_{j}=E_{j}+\epsilon_{jkl}\dot{x}_{k}B_{l},$ (2.4)
which is the Lorentz force equation for a particle of unit charge. Taking eq.(2.4) as definition of the field E, we see $[x_{j},E_{k}]=0$. Rewriting eq.(2.3) as
$B_{l}=-\frac{im^{2}}{2\hbar}\epsilon_{ljk}[\dot{x}_{j},\dot{x}_{k}]$ (2.5)
and applying Jacobi identity involving velocities, we get
$[\dot{x}_{l},B_{l}]=0$ (2.6)
which leads to
$\partial_{i}B_{i}=0$ (2.7)
By taking total time derivative of the definition of $B$ given in eq.(2.5), the second Maxwell’s equation is obtained as,
$\frac{\partial B_{l}}{\partial t}=\epsilon_{ljk}\frac{\partial E_{j}}{\partial x_{k}}$ (2.8)
The other two Maxwell’s equations define the charge and current density.
In [32], Feynman’s approach is generalised to relativistic case and extended to non-abelian gauge theory, generalising the results of [31]. In [32], a particle carrying isospin $I^{a}$ satisfying,
$m[x_{\mu},\dot{x}_{\nu}]=-i\hbar\eta_{\mu\nu},~{}~{}\eta_{\mu\nu}={\rm diag}(1,-1,-1,-1)$ (2.9)
$[I^{a},I^{b}]=i\hbar f^{abc}I^{c};~{}~{}~{}[x^{\mu},I^{a}]=0$ (2.10)
and the Wong’s equation
$\dot{I}^{a}-f^{abc}A^{b}_{\mu}\dot{x}^{\mu}I^{c}=0$ (2.11)
is considered. Here one follows the same prescription as in [30], starting by differentiating eq.(2.9) with respect to $\tau$ and define,
$F^{\mu\nu}=\frac{m}{i\hbar}[x^{\mu},F^{\nu}]=-\frac{m^{2}}{i\hbar}[\dot{x}^{\mu},\dot{x}^{\nu}]$ (2.12)
where one used $F^{\nu}=m\ddot{x}^{\nu}$. By using Jacobi identity it is showed that $F_{\mu\nu}$ is a function of $x^{\mu}$ alone which allows the integration of the above expression resulting in
$F^{\mu}(x,\dot{x},t)=G^{a\mu}(x)I^{a}+<F^{a\mu}_{\nu}I^{a}\dot{x}^{\nu}>$ (2.13)
Using Wong’s equation, the commutator of four-velocities with Lie algebra valued functions is calculated. This also provide us with the definition of gauge covariant derivatives. Then by using the Jacobi identity involving velocities and the definition of $F^{\mu\nu}$ given in eq.(2.12), one obtains
$D_{\mu}F_{\nu\rho}+D_{\nu}F_{\rho\mu}+D_{\rho}F_{\mu\nu}=0$ (2.14)
Also by using the commutation relation $[\dot{x}_{\mu},I^{a}]$, obtained using Wong’s equation and Jacobi identities one finds the explicit expression for $F^{a}_{\mu\nu}$ in terms of gauge fields as $F^{a}_{\mu\nu}=\partial_{\mu}A^{a}_{\nu}-\partial_{\nu}A^{a}_{\mu}-ef^{abc}A^{b}_{\mu}A^{c}_{\nu}$. In this approach, the homogeneous Yang-Mills equations and force equation are obtained. In our study, we generalise this approach to the $SU(N)$ gauge theory in the $\kappa$-deformed space-time.
3 Wong’s equation in $\kappa$-space-time
The generators $I^{a}$, $a=1,2,..n^{2}-1$ of $su(n)$ algebra satisfy the relation
$[I^{a},I^{b}]=i\hbar f^{abc}I^{c}.$ (3.1)
For deriving the Wong’s equation in $\kappa$-deformed space-time, we start from,
$\frac{dI^{a}}{dt}=\frac{i}{\hbar}[\hat{H},I^{a}].$ (3.2)
where $\hat{H}$ is the Dirac Hamiltonian in $\kappa$-deformed space-time. The Dirac equation in $\kappa$-deformed space is[26]
$(i\gamma^{\mu}D_{\mu}-\frac{mc}{\hbar})\psi=0,$ (3.3)
where $\gamma^{\mu}$ are the Dirac matrices satisfying the relations $(\gamma^{0})^{2}=\mathbb{I};~{}~{}~{}\gamma^{0}\gamma^{i}=\alpha^{i};~{}~{}[\gamma^{\mu},\gamma^{\nu}]_{+}=2\eta^{\mu\nu}$.
In our study, non-commutative coordinates are expressed in terms of commutative variables [23] as,
$\hat{x}_{0}=x_{0};~{}~{}~{}\hat{x}_{i}=x_{i}(1-\frac{a}{\hbar}p_{0})$ (3.4)
where $x_{0}$, $x_{i}$ are the coordinates of the commutative space-time and $p_{0}$ is the conjugate momentum corresponding to $x_{0}$. The components of the Dirac derivatives appearing in the above are [26] are
$D_{0}=\partial_{0}+\frac{ia}{2}\nabla^{2};~{}~{}~{}D_{i}=\partial_{i},$ (3.5)
where $\nabla^{2}=\partial_{i}\partial_{i}$. Note that $\partial_{i}$ and $\partial_{0}$ are the derivatives corresponding to the commutative coordinates. Here we have considered corrections valid only upto first order in the deformation parameter, $a$. Substituting eq.(3.5) in eq.(3.3) gives
$\Big{(}i\gamma^{0}(\partial_{0}+\frac{ia}{2}\nabla^{2})+i\gamma^{i}\partial_{i}-\frac{mc}{\hbar}\Big{)}\psi=0.$ (3.6)
Using $p_{0}=i\hbar\partial_{0}$ and $p_{i}=i\hbar\partial_{i}$ and applying the minimal coupling prescription, i.e.,
$p_{0}\rightarrow p_{0}-eA_{0},~{}~{}p_{i}\rightarrow p_{i}-eA_{i},$ (3.7)
where $A_{\mu}$, $\mu=0,1,2,3$ are the components of the four-vector potential, and simplifying we get $\kappa$-deformed Dirac equation in presence of external non-abelian gauge field as,
$i\hbar\frac{\partial\psi}{\partial t}=\Big{(}-c\alpha^{i}(p_{i}-eA_{i})+ecA_{0}-\frac{ac}{2\hbar}(p_{i}-eA_{i})^{2}+mc^{2}\beta\Big{)}\psi,$ (3.8)
where $\gamma_{0}$ is re-expressed as $\beta$. From the above equation, we read off the $\kappa$-deformed Dirac Hamiltonian to be
$\hat{H}=-c\alpha^{i}(p_{i}-eA_{i})+ecA_{0}-\frac{ac}{2\hbar}(p_{i}-eA_{i})^{2}+mc^{2}\beta.$ (3.9)
Note that the $a$-dependent term in the above modified Hamiltonian is coming from the Dirac derivative. We now use this modified Dirac Hamiltonian for finding the Wong’s equation in $\kappa$-deformed space-time. Substituting Hamiltonian obtained in eq.(3.9) in eq.(3.2) and using $A_{\mu}=A^{a}_{\mu}I^{a}$ gives
$\frac{dI^{a}}{dt}$ $=\frac{i}{\hbar}\bigg{[}-c\alpha^{i}(p_{i}-eA^{b}_{i}I^{b})+ecA^{b}_{0}I^{b}-\frac{ac}{2\hbar}(p_{i}-eA^{b}_{i}I^{b})^{2}+mc^{2}\beta,~{}I^{a}\bigg{]}$
$=\frac{i}{\hbar}\bigg{[}ce\alpha^{i}A^{b}_{i}[I^{b},I^{a}]+ecA^{b}_{0}[I^{b},I^{a}]+\frac{ace}{2\hbar}\Big{(}(p_{i}-eA_{i})A^{b}_{i}[I^{b},I^{a}]+A^{b}_{i}[I^{b},I^{a}](p_{i}-eA_{i})\Big{)}\bigg{]}.$ (3.10)
Using the ordering prescription where $x^{\mu}$ is always written to the left of $p^{\mu}$ for the third term in the above expression, we re-write the above equation as
$\frac{dI^{a}}{dt}$ $=ec\alpha^{i}f^{abc}A^{b}_{i}I^{c}+ecf^{abc}A^{b}_{0}I^{c}+\frac{ace}{\hbar}f^{abc}p_{i}A^{b}_{i}I^{c}-\frac{ace^{2}}{2\hbar}f^{abc}A^{b}_{i}A^{d}_{i}[I^{d}I^{c}+I^{c}I^{d}]$ (3.11)
Using Hamiltonian in eq.(3.9) we now evaluate
$cD_{0}\hat{x}_{j}=\frac{i}{\hbar}[\hat{H},\hat{x}_{j}]$ (3.12)
Using the realisation for $D_{0}$ and $\hat{x}_{j}$ given in eq.(3.5) and eq.(3.4), respectively, we find
$D_{0}\hat{x}_{j}=(\partial_{0}+\frac{ia}{2}\nabla^{2})(x_{j}-\frac{a}{\hbar}x_{j}p_{0}),~{}~{}=\partial_{0}x_{j}(1-\frac{ap_{0}}{\hbar})-\frac{a}{\hbar c}x_{j}\frac{dp_{0}}{dt}~{}~{}\cong\partial_{0}x_{j}e^{-\frac{ap_{0}}{\hbar}}-\frac{a}{\hbar c}x_{j}\frac{dp_{0}}{dt},$ (3.13)
Here we are considering corrections valid only upto first order in deformation parameter $a$ and hence approximated $(1-\frac{ap_{0}}{\hbar})=e^{-\frac{ap_{0}}{\hbar}}$. The RHS of eq.(3.12) is calculated using Hamiltonian in eq.(3.9) as
$\frac{i}{\hbar}[\hat{H},\hat{x}_{j}]=c\alpha_{j}-\frac{ac}{\hbar}(p_{j}-eA_{j})-\frac{a}{\hbar}c\alpha_{j}p_{0}.$ (3.14)
Equating these two equations, we get $c\alpha_{i}=c\partial_{0}x_{i}-\frac{a}{\hbar}x_{i}\frac{dp_{0}}{dt}+\frac{ac}{\hbar}(p_{i}-eA_{i})$. Substituting this in eq.(3.11), we find
$\frac{dI^{a}}{d\tau}=e\dot{x}^{\mu}f^{abc}A^{b}_{\mu}I^{c}-\frac{ace^{2}}{2\hbar}f^{abc}A^{b}_{i}A^{d}_{i}(I^{c}I^{d}-I^{d}I^{c})$ (3.15)
where we have replaced $t$ with $\tau$ as we are in the relativistic case [32]. This is the $\kappa$-modified Wong’s equation. Note that the non-commutative terms coming in the modified Wong’s equation through the Dirac derivative (see eq.(3.5)) alone and not from the realisation in eq.(3.4). Here we have also used the ordering prescription for x and p.
4 Yang-Mills equation in the $\kappa$-deformed space-time
In $\kappa$-deformed space-time, the position coordinates obey the commutation relation given by eq.(1.1). Differentiating eq.(1.1) with respect to the parameter $\tau$ gives
$[\dot{\hat{x}}_{\mu},\hat{x}_{\nu}]+[\hat{x}_{\mu},\dot{\hat{x}}_{\nu}]=i(a_{\mu}\dot{\hat{x}}_{\nu}-a_{\nu}\dot{\hat{x}}_{\mu})$ (4.1)
Since eq.(4.1) is symmetric under the exchange of indices $\mu$ and $\nu$, we introduce a symmetric tensor $\hat{\eta}_{\mu\nu}$ as
$[\hat{x}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar}{m}\hat{\eta}_{\mu\nu}.$ (4.2)
Differentiating the above equation with respect to $\tau$ gives
$[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]+[\hat{x}_{\mu},\ddot{\hat{x}}_{\nu}]=-\frac{i\hbar}{m}\frac{d\hat{\eta}_{\mu\nu}}{d\tau}$ (4.3)
Note that the first term in LHS of eq.(4.3) is antisymmetric under exchange of indices $\mu$ and $\nu$ and this allows us to introduce an anti-symmetric tensor $\hat{F}_{\mu\nu}$, through the definition,
$m[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar e}{m}\hat{F}_{\mu\nu}+\hat{S}_{\mu\nu}$ (4.4)
where $\hat{S}_{\mu\nu}=-\frac{i\hbar}{m}\frac{d\hat{\eta}_{\mu\nu}}{d\tau}=-\frac{i\hbar}{m}\frac{d\delta\hat{\eta}_{\mu\nu}}{d\tau}$ as the commutative metric is constant in $\tau$. Thus $\hat{S}_{\mu\nu}$ has no commutative part. Next consider the commutation relation
$[\hat{x}_{0},I^{a}]=0.$ (4.5)
Differentiating eq.(4.5)with respect to $\tau$, we get
$[\dot{\hat{x}}_{0},I^{a}]+[\hat{x}_{0},\dot{I}^{a}]=0.$ (4.6)
Substituting Wong’s equation obtained in eq.(3.15) for $\dot{I}^{a}$ in the second term and rearranging the above expression gives
$[\dot{\hat{x}}_{0},I^{a}]=\frac{i\hbar e}{m}f^{abc}A^{b}_{0}I^{c}.$ (4.7)
Here as $[x_{0},I^{a}]=0$, the $a$-dependent term in Wong’s equation does not contribute to the above expression. Next we find the commutation relation between $\dot{\hat{x}}_{0}$ and Lie algebra valued functions of $\kappa$-deformed coordinates. For this we start with an arbitrary function $\phi^{a}(\hat{x}_{\mu})$ and consider its Taylor series expansion given by $\phi^{a}(\hat{x}_{\mu})=\phi^{a}(x_{\mu}-\frac{a}{\hbar}x_{\mu}p_{0})=\phi(x_{\mu})-\frac{a}{\hbar}x_{\nu}\partial^{\nu}\phi^{a}p_{0}$. Using this and the realisation of non-commutative coordinates in terms of commutative coordinates (see eq.(3.4) and eq.(4.7)), we get
$[\dot{\hat{x}}_{0},\phi^{a}(\hat{x}_{\mu})I^{a}]=\frac{i\hbar}{m}(\delta^{ac}\partial_{0}-ef^{abc}A^{b}_{0}){\phi}^{c}I^{a}=\frac{i\hbar}{m}\tilde{D}^{ac}_{0}\phi^{c}I^{a}$ (4.8)
where $\tilde{D}^{ac}_{0}\phi^{c}=(\delta^{ac}\partial_{0}-ef^{abc}A^{b}_{0})\phi^{c}$ is the gauge covariant derivative in $\kappa$-deformed space-time. Note that $\tilde{D}_{0}$ does not get any $a$ dependent correction. Similarly, using
$[\dot{\hat{x}}_{i},I^{a}]=\frac{i\hbar e}{m}f^{abc}A^{b}_{i}I^{c}-\frac{iae}{m}f^{abc}A^{b}_{i}I^{c}p_{0},$ (4.9)
${\rm~{}we~{}find~{}~{}}[\dot{\hat{x}}_{i},\phi^{a}(\hat{x}_{\mu})I^{a}]=(1-\frac{ap_{0}}{\hbar})\frac{i\hbar}{m}\Big{(}\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}\Big{)}{\phi}^{c}I^{a}=\frac{i\hbar}{m}\tilde{D}^{ac}_{i}\phi^{c}I^{a}$ (4.10)
where $\tilde{D}_{i}^{ac}\phi^{c}=e^{-\frac{ap_{0}}{\hbar}}(\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}){\phi}^{c}$ is the spatial part of the gauge covariant derivative in $\kappa$-deformed space-time. Here all the $a$-dependent terms are coming from the realisation of $\hat{x}_{i}$ (see eq.(3.4)). As we are considering terms only upto first order in the deformation parameter $a$, $1-\frac{ap_{0}}{\hbar}$ is taken as $e^{-\frac{ap_{0}}{\hbar}}$. Thus we obtain all the components of $\kappa$-deformed gauge covariant derivative. Using the Jacobi identity $[\dot{\hat{x}}_{\mu},[\dot{\hat{x}}_{\nu},\dot{\hat{x}}_{\rho}]]+[\dot{\hat{x}}_{\nu},[\dot{\hat{x}}_{\rho},\dot{\hat{x}}_{\mu}]]+[\dot{\hat{x}}_{\rho},[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]]=0$ and substituting eq.(4.4) in the above, we find that the $\hat{F}_{\mu\nu}$ satisfies
$\epsilon^{0ijk}(\tilde{D}_{i}\hat{F}_{jk}+\tilde{D}_{j}\hat{F}_{ki}+\tilde{D}_{k}\hat{F}_{ij})=0,$ (4.11)
$\epsilon^{0ijk}(\tilde{D}_{0}\hat{F}_{jk}+2\tilde{D}_{j}\hat{F}_{k0})=0.$ (4.12)
where we used $[\dot{\hat{x}}^{\mu},\frac{d\delta\hat{\eta}^{\nu\rho}}{d\tau}]=0$ as the $\kappa$-deformed corrections to the metric is known to be independent of $\hat{x}^{\mu}$ and depend only on $p^{\mu}$ [34, 35, 36]. It remains to be shown that $\hat{F}_{ij}$ and $\hat{F}_{0i}$ are indeed $\kappa$-deformed Yang-Mills field strength. To show this, we start from the commutation relation between $\dot{\hat{x}}_{0}$ and $I^{a}$ given in eq.(4.7). Taking commutator of eq.(4.7) with $\dot{\hat{x}}_{i}$ gives,
$0=[\dot{\hat{x}}_{i},[\dot{\hat{x}}_{0},I^{a}]]-\frac{i\hbar e}{m}f^{abc}[\dot{\hat{x}}_{i},A^{b}_{0}I^{c}]$ (4.13)
Using the Jacobi idenity to re-express the first term as $[\dot{\hat{x}}_{0},[\dot{\hat{x}}_{i},I^{a}]]+[I^{a},[\dot{\hat{x}}_{0},\dot{\hat{x}}_{i}]]$ and eq.(3.4), the above equation becomes
00 0 $=$ $[\dot{x}_{0},[\dot{\hat{x}}_{i},I^{a}]]-\frac{i\hbar e}{m^{2}}[I^{a},\hat{F}^{b}_{0i}I^{b}]-\frac{i\hbar e}{m}f^{abc}\Big{(}[\dot{x}_{i}-\frac{a}{\hbar}\dot{x}_{i}p_{0},A^{b}_{0}I^{c}]\Big{)}$ (4.14)
In writing the second term in the above equation, we have used the eq.(4.4). Using eq.(4.9) and eq.(3.1), we get the expression for $\hat{F}^{a}_{0i}$ as
$\hat{F}^{a}_{0i}=(\partial_{0}A^{a}_{i}-\partial_{i}A^{a}_{0}-ef^{abc}A^{b}_{0}A^{c}_{i})(1-\frac{ap_{0}}{\hbar})=F^{a}_{0i}e^{-\frac{ap_{0}}{\hbar}}$ (4.15)
As earlier we have taken $1-\frac{ap_{0}}{\hbar}$ to be $e^{-\frac{ap_{0}}{\hbar}}$. Here $F^{a}_{0i}=\partial_{0}A^{a}_{i}-\partial_{i}A^{a}_{0}-ef^{abc}A^{b}_{0}A^{c}_{i}$, is the commutative “Electric field”.
Next by taking commutator of eq.(4.9) with $\dot{\hat{x}}_{j}$, we get
$0=[\dot{\hat{x}}_{j},[\dot{\hat{x}}_{i},I^{a}]]-\frac{i\hbar e}{m}f^{abc}([\dot{\hat{x}}_{j},A^{b}_{i}]I^{c}+A^{b}_{i}[\dot{\hat{x}}_{j},I^{c}])+\frac{iae}{m}f^{abc}([\dot{\hat{x}}_{j},A^{b}_{i}I^{c}p_{0}]).$ (4.16)
As earlier applying the Jacobi identity in the first term and using eq.(4.4), we find
$0=$ $~{}[\dot{\hat{x}}_{i},[\dot{\hat{x}}_{j},I^{a}]]-\frac{i\hbar e}{m^{2}}\hat{F}^{b}_{ij}[I^{a},I^{b}]-\frac{i\hbar e}{m}f^{abc}([\dot{\hat{x}}_{j},A^{b}_{i}]I^{c}+A^{b}_{i}[\dot{\hat{x}}_{j},I^{c}])+\frac{iae}{m}f^{abc}([\dot{\hat{x}}_{j},A^{b}_{i}I^{c}p_{0}])$ (4.17)
Next using eq.(4.9) and eq.(3.1) in the right hand side of the above equation and simplifying, we get
$\hat{F}_{ij}^{a}=(\partial_{i}A^{a}_{j}-\partial_{j}A^{a}_{i}-ef^{abc}A^{b}_{i}A^{c}_{j})(1-\frac{2ap_{0}}{\hbar})=F^{a}_{ij}e^{-\frac{2ap_{0}}{\hbar}}$ (4.18)
where NC correction is approximated as the exponential and $F^{a}_{ij}=\partial_{i}A^{a}_{j}-\partial_{j}A^{a}_{i}-ef^{abc}A^{b}_{i}A^{c}_{j}$ is the commutative “magnetic” field.
We observe that the second rank anti-symmetric tensor introduced in eq.(4.4) is indeed the $\kappa$-deformed Yang-Mills field strength. Note that the $a$-dependent term of the Yang-Mills field strength is coming due to the non-commutativity of the coordinates. This form of $F_{\mu\nu}$ is compatible with
$([\tilde{D}_{0},\tilde{D}_{i}]\phi)^{a}=-ef^{abc}F^{b}_{0i}\phi^{c}(1-\frac{ap_{0}}{\hbar})=-ef^{abc}F^{b}_{0i}\phi^{c}e^{-{\frac{ap_{0}}{\hbar}}}$ (4.19)
valid upto first order in $a$, similarly
$([\tilde{D}_{i},\tilde{D}_{j}]\phi)^{a}=-ef^{abc}F^{b}_{ij}\phi^{c}(1-\frac{2ap_{0}}{\hbar})=-ef^{abc}F^{b}_{ij}\phi^{c}e^{-{\frac{2ap_{0}}{\hbar}}}$ (4.20)
We observe that the commutators of components of the $\kappa$-deformed gauge covariant derivatives exactly matches with the deformed Yang-Mills field strength we derived in $\kappa$-space-time (see eq.(4.15) and eq.(4.18)). We also note that the $a$-dependent corrections to $\hat{F}^{a}_{\mu\nu}$ are coming solely from the non-commutativity of the coordinates, and not from the $a$-dependent corrections of Wong’s equation. This shows that eq.(4.11) and eq.(4.12) are the Yang-Mills equations.
5 Yang-Mills Lagrangian in $\kappa$-space-time
In the non-commutative space-time, we have found that $\hat{F}_{ij}^{a}=F^{a}_{ij}(1-\frac{2ap_{0}}{\hbar})$ and $\hat{F}^{a}_{0i}=F^{a}_{0i}(1-\frac{ap_{0}}{\hbar})$, where $F_{ij}$ and $F_{0i}$ are the commutative field strengths. Using these, we find the Lagrangian as,
$\hat{L}=-\frac{1}{2}(1-\frac{ap_{0}}{\hbar})^{2}F^{a}_{0i}F^{a0i}-\frac{1}{4}(1-2\frac{ap_{0}}{\hbar})^{2}F^{a}_{ij}F^{aij},$ (5.1)
is invariant under the usual SU(N) gauge transformations. We have already derived two of the Yang-Mills equations (see eq.(4.11) and eq.(4.12)). The remaining two equations, follow as the Euler-Lagrange equation from the above Lagrangian, and they are
$(1-\frac{2ap_{0}}{\hbar})(D_{i}F^{0i})^{a}=0,~{}~{}(D_{0}F^{i0})^{a}+(1-\frac{2ap_{0}}{\hbar})(D_{j}F^{ij})^{a}=0$ (5.2)
In the limit $a\rightarrow 0$, above equations reduces to the covariant commutative result. It is clear that under the gauge transformation of the vector field, $A^{a}_{\mu}\rightarrow A^{a}_{\mu}+(D_{\mu}\alpha(x))^{a}$, where the covariant derivative is $(D_{\mu}\alpha(x))^{a}=(\delta^{ac}\partial_{\mu}-ef^{abc}A^{b}_{\mu}){\alpha}^{c}(x)$, the Lagrangian given in eq.(5.1) is invariant, showing the invariance of $\kappa$-deformed Yang-Mills field theory under usual SU(N) transformation.
6 Lorentz force equation in presence of SU(N)gauge fields
To derive the force experienced by an isospin carrying particle in presence of Yang-Mills field in $\kappa$-space-time, we start with $[\hat{x}_{i},\hat{F}_{0}]$ where $\hat{F}_{0}$ is the zeroth-component of the 4-vector force. We first do a Taylor series expansion of $\hat{F}_{0}$, considering it to be a function of $\hat{x}_{0}$ and $\hat{x}_{i}$. We then use this Taylor series expansion of $\hat{F}_{0}$, $\hat{F}_{0}=F_{0}-\frac{a}{\hbar}(x_{j}\partial^{j}F_{0}p_{0}+\dot{x_{j}}\frac{\partial F_{0}}{\partial\dot{x}_{j}}p_{0})$ in the above equation. For $\hat{x}_{i}$ we use the realisation given in eq.(3.4). Thus we find
$[\hat{x}_{i},\hat{F}_{0}]=[x_{i}-\frac{a}{\hbar}x_{i}p_{0},~{}F_{0}-\frac{a}{\hbar}(x_{j}\partial^{j}F_{0}p_{0}+\dot{x_{j}}\frac{\partial F_{0}}{\partial\dot{x}_{j}}p_{0})]$ (6.1)
Keeping corrections valid only upto first order in the deformation parameter $a$, we find
$[\hat{x}_{i},\hat{F}_{0}]=\frac{i\hbar e}{m}F_{i0}-\frac{a}{\hbar}\Big{(}{\frac{i\hbar e}{m}}F_{i0}p_{0}-\frac{i\hbar}{m}x_{j}\frac{\partial(\partial^{j}F_{0})}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\frac{\partial F_{0}}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{j}\frac{\partial(\frac{\partial F_{0}}{\partial\dot{x}_{j}})}{\partial\dot{x}^{i}}p_{0}\Big{)}$ (6.2)
Consider $\hat{F}_{0}=F_{0}+a\delta F_{0}$ where $\delta F_{0}$ is the modified part of force in the $\kappa$- deformed space-time and then we find the commutator,
$[\hat{x}_{i},\hat{F}_{0}]=[x_{i}-\frac{a}{\hbar}x_{i}p_{0},~{}F_{0}+a\delta F_{0}]$ (6.3)
From eq.(6.2) and eq.(6.3), we equate the commutative and non-commutative part separately. Equating commutative part gives, $\frac{i\hbar e}{m}F_{i0}=-\frac{i\hbar}{m}\frac{\partial F_{0}}{\partial\dot{x}^{i}}$. which on integration gives,
$F_{0}=eF_{0i}\dot{x}^{i}+G_{0}(x)$ (6.4)
where $G_{0}(x)$, is a function of x. Now equating the non-commutative part of the equations eq.(6.2) and eq.(6.3) gives,
$-\frac{i\hbar}{m}\frac{\partial\delta F_{0}}{\partial\dot{x}^{i}}=-\frac{1}{\hbar}\bigg{[}{\frac{i\hbar e}{m}}F_{i0}p_{0}-\frac{i\hbar}{m}x_{j}\frac{\partial(\partial^{j}F_{0})}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{j}\frac{\partial(\frac{\partial F_{0}}{\partial\dot{x}_{j}})}{\partial\dot{x}^{i}}p_{0}\bigg{]}$ (6.5)
By substituting $F_{0}$ from the commutative result (eq.(6.4)) and integrating, we get
$\delta F_{0}=\frac{1}{\hbar}[eF_{i0}P_{0}-ex_{j}\partial^{j}F_{0i}p_{0}]\dot{x}^{i}+\delta{G}_{0}$ (6.6)
Using eq.(6.4) and eq.(6.6), we find
$\hat{F}_{0}=eF_{0i}\dot{x}^{i}(1-\frac{a}{\hbar}p_{0})-\frac{ae}{\hbar}x_{j}\partial^{j}F_{0i}\dot{x}^{i}p_{0}+\hat{G}_{0}$ (6.7)
where $\hat{G}_{0}=G_{0}+a\delta{G}_{0}$. Similarly, for deriving the expression for $\hat{F}_{i}$, consider
$[\hat{x}_{i},\hat{F}_{j}]=\bigg{[}x_{i}-\frac{a}{\hbar}x_{i}p_{0},~{}F_{j}-\frac{a}{\hbar}(x_{k}\partial^{k}F_{j}p_{0}+\dot{x}_{k}\frac{\partial F_{j}}{\partial\dot{x}_{k}}p_{0})\bigg{]}$ (6.8)
Here we have used the realisation of non-commutative coordinates in terms of commutative coordinates and used Taylor series expansion of $\hat{F}_{i}$, valid upto first order in $a$. Using $[x_{i},F_{j}]=\frac{i\hbar e}{m}F_{ij}$ we get,
$[\hat{x}_{i},\hat{F}_{j}]=\frac{i\hbar e}{m}F_{ij}-\frac{a}{\hbar}\Big{(}{\frac{i\hbar e}{m}}F_{ij}p_{0}-\frac{i\hbar}{m}x_{k}\frac{\partial(\partial^{k}F_{j})}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\frac{\partial F_{j}}{\partial\dot{x}^{i}}p_{0}-\frac{i\hbar}{m}\dot{x}_{k}\frac{\partial(\frac{\partial F_{j}}{\partial\dot{x}_{k}})}{\partial\dot{x}^{i}}p_{0}\Big{)}$ (6.9)
Taking $\hat{F}_{j}={F}_{j}+a\delta\bar{F}_{j}$, we find
$[\hat{x}_{i},\hat{F}_{j}]=[x_{i}-\frac{a}{\hbar}x_{i}p_{0},~{}F_{j}+a\delta\bar{F}_{j}]$ (6.10)
and equating the commutative part of eq.(6.9) and eq.(6.10) gives, $\frac{i\hbar e}{m}F_{ij}=-\frac{i\hbar}{m}\frac{\partial F_{j}}{\partial\dot{x}^{i}}$. which on integration gives
$F_{j}=eF_{ji}\dot{x}^{i}+\bar{G}_{j}(x)$ (6.11)
From the non-commutative part of eq.(6.9) and eq.(6.10), we get
$-\frac{i\hbar}{m}\frac{\partial\delta\bar{F}_{j}}{\partial\dot{x}^{i}}=\frac{i}{m}\bigg{[}x_{k}\frac{\partial(\partial^{k}F_{j})}{\partial\dot{x}^{i}}p_{0}+\frac{\partial F_{j}}{\partial\dot{x}^{i}}p_{0}+\dot{x}_{k}\frac{\partial(\frac{\partial F_{j}}{\partial\dot{x}_{k}})}{\partial\dot{x}^{i}}p_{0}\bigg{]}$ (6.12)
Using the commutative expression for $F_{j}=eF_{ji}\dot{x}^{i}+eF_{j0}\dot{x}^{0}+G_{j}$ [32], we get, $\frac{\partial\delta\bar{F}_{j}}{\partial\dot{x}^{i}}=-\frac{1}{\hbar}(ex_{k}\partial^{k}F_{ji}p_{0}+eF_{ji}p_{0})$. On integrating the above equation gives,
$\delta\bar{F}_{j}=-\frac{1}{\hbar}(ex_{k}\partial^{k}F_{ji}p_{0}\dot{x}^{i}+eF_{ji}p_{0}\dot{x}^{i})+\delta\bar{G}_{j}$ (6.13)
Thus, we get
$\hat{F}_{i}=eF_{ij}\dot{x}^{j}(1-\frac{ap_{0}}{\hbar})+\hat{G^{\prime}_{i}}(x)-\frac{a}{\hbar}ex_{k}\partial^{k}F_{ij}\dot{x}^{j}p_{0}$ (6.14)
Here we are considering terms only upto first order in $a$ and thus $\hat{G^{\prime}}_{i}=\bar{G}_{i}+a\delta\bar{G}_{i}$. For getting the complete expression of $\hat{F}_{i}$, we consider $[\hat{x}_{0},\hat{F}_{i}]$ also, i.e.,
$[\hat{x}_{0},\hat{F}_{i}]$ $=[x_{0},F_{i}-\frac{a}{\hbar}(x_{k}\partial^{k}F_{i}p_{0}+\dot{x}_{k}\frac{\partial F_{i}}{\partial\dot{x}_{k}}p_{0})]$
$=\frac{i\hbar e}{m}F_{0i}+\frac{a}{\hbar}\Big{(}\frac{i\hbar}{m}x_{k}\frac{\partial(\partial^{k}F_{i})}{\partial\dot{x}_{0}}p_{0}+i\hbar x_{k}\partial^{k}F_{i}+\frac{i\hbar}{m}\dot{x}_{k}\frac{\partial(\frac{\partial F_{i}}{\partial\dot{x}_{k}})}{\partial\dot{x}_{0}}p_{0}+i\hbar\dot{x}_{k}\frac{\partial F_{i}}{\partial\dot{x}_{k}}\Big{)}$ (6.15)
As earlier, we consider $\hat{F}_{i}=F_{i}+a\delta F_{i}$, and find
$[\hat{x}_{0},\hat{F}_{i}]=[x_{0},F_{i}+a\delta F_{i}]=-\frac{i\hbar}{m}\frac{\partial F_{i}}{\partial\dot{x}_{0}}-\frac{i\hbar a}{m}\frac{\partial\delta F_{i}}{\partial\dot{x}_{0}}$ (6.16)
On equating commutative and non-commutative part, we get
$F_{i}=eF_{i0}\dot{x}_{0}+G_{i}$ (6.17)
$\delta F_{i}$ $=-\frac{1}{\hbar}\Big{(}ex_{k}\partial^{k}F_{i0}p_{0}\dot{x}_{0}+mex_{k}\partial^{k}F_{ij}\dot{x}^{j}\dot{x}_{0}+mex_{k}\partial^{k}F_{i0}\frac{\dot{x}_{0}^{2}}{2}+mx_{k}\partial^{k}G_{i}\dot{x}_{0}+$
$~{}~{}~{}~{}me\dot{x}_{k}F^{k}_{i}\dot{x}_{0}\Big{)}+\delta{G}_{i}$ (6.18)
Combining eq.(6.14) eq.(6.17) and eq.(6) we get the expression for force in $\kappa$-deformed space-time as,
$\hat{F}_{i}$ $=eF_{i\mu}\dot{x}^{\mu}+\hat{G_{i}}(x)-\frac{a}{\hbar}\Big{(}eF_{ij}\dot{x}^{j}p_{0}+ex_{k}\partial^{k}F_{i\mu}\dot{x}^{\mu}p_{0}$
$\quad+mex_{k}\partial^{k}F_{ij}\dot{x}^{j}\dot{x}_{0}+mex_{k}\partial^{k}F_{i0}\frac{\dot{x}_{0}^{2}}{2}+mx_{k}\partial^{k}G_{i}\dot{x}_{0}+me\dot{x}_{k}F^{k}_{i}\dot{x}_{0}\Big{)}.$ (6.19)
Eq.(6.7) and eq.(6) gives the components of force experienced by an isospin carrying particle in presence of Yang-Mills field in $\kappa$-deformed space-time.
From eq.(6.7), we can write
$\hat{G}_{0}=\hat{F}_{0}-eF_{0i}\dot{x}_{i}(1-\frac{a}{\hbar}p_{0})+\frac{ae}{\hbar}x_{j}\partial^{j}F_{0i}\dot{x}^{i}p_{0}$ (6.20)
Similarly from eq.(6),
$\hat{G}_{i}$ $=\hat{F}_{i}-eF_{i\mu}\dot{x}^{\mu}+\frac{a}{\hbar}\Big{(}eF_{ij}\dot{x}^{j}p_{0}+ex_{k}\partial^{k}F_{i\mu}\dot{x}^{\mu}p_{0}$
$\quad+mex_{k}\partial^{k}F_{ij}\dot{x}^{j}\dot{x}_{0}+mex_{k}\partial^{k}F_{i0}\frac{\dot{x}_{0}^{2}}{2}+mx_{k}\partial^{k}G_{i}\dot{x}_{0}+me\dot{x}_{k}F^{k}_{i}\dot{x}_{0}\Big{)}.$ (6.21)
Using these, we calculate $\tilde{D}_{0}\hat{G}_{i}-\tilde{D}_{i}\hat{G}_{0}$ and $\tilde{D}_{i}\hat{G}_{j}-\tilde{D}_{j}\hat{G}_{i}$ and find that both have non-zero non-commutative correction.
7 Conclusion
We have constructed Yang-Mills theory in $\kappa$-deformed space-time, by generalising Feynman’s approach. For this, we have first constructed the Wong’s equation in $\kappa$-deformed space-time. We have obtained the force equation for a particle having isospin in presence of non-abelian gauge field. We have also constructed the Lagrangian for $\kappa$-deformed Yang-Mills theory and showed its invariance under SU(N) gauge transformations. Thus we see that the $\kappa$-deformed Yang-Mills theory obtained has the same gauge invariance as its commutative counter part. In our study, we have considered corrections valid only up to first order in the deformation parameter $a$.
In our study, by considering the anti-symmetry property of commutators between velocities, we have introduced a second-rank, anti-symmetric tensor, which was later found to be the Yang-Mills field strength tensor in $\kappa$-deformed space-time. Using the Jacobi identity repeatedly, we have showed that this second rank, anti-symmetric tensor obey homogeneous Yang-Mills field equation. In evaluating the Jacobi identity, we need commutation relation between Lie algebra valued functions and $\kappa$-deformed velocities, which in turn leads to the commutation relation between generators of Lie algebra and the velocities. This is obtained by using Wong’s equation in $\kappa$-space-time, which we constructed using the $\kappa$-deformed Dirac Hamiltonian. Note that the $\kappa$-deformed Dirac equation is constructed [26] using the Dirac derivatives [23] which are written in terms of derivatives with respect to the commutative space-time coordinates and the deformation parameter. Thus the $\kappa$-Dirac equation in momentum space will depend only on commutative momentum $p_{0}$, $p_{i}$. Therefore, the minimal coupling method used to bring in the gauge interaction allows us to introduce only gauge field living in the commutative space-time. Here if by brute force we use $\hat{A}^{\mu}(\hat{x})$, eq.(4.7) will have $a$-dependent term in the RHS while LHS is independent of any, leading to inconsistency. Thus we use $A(x)$ in implementing minimal coupling in eq.(3.8). Note that in the $\kappa$-space-time, it is the time coordinate that does not commute with space-coordinates while space coordinates commute among themselves. This should be contrasted with the Moyal space-time and thus the no-go theorem obtained in Moyal space-time [37] is not applicable to our study. By repeated use of Jacobi identity, then we have derived two of the homogeneous Yang-Mills field equation in $\kappa$-deformed space-time.
We have obtained the explicit expression for $\hat{F}_{0i}$ and $\hat{F}_{ij}$ and found their $a$-dependent corrections. The correction to $\hat{F}_{0i}$ goes like $e^{-\frac{ap_{0}}{\hbar}}$ whereas the non-commutative correction to $\hat{F}_{ij}$ goes as $e^{-\frac{2ap_{0}}{\hbar}}$. We have also obtained the explicit expression for the gauge covariant derivative $\tilde{D}_{0}$ and $\tilde{D}_{i}$ and observed that $\tilde{D}_{0}$ does not get any $a$-dependent correction while $\tilde{D}_{i}$ does have $a$-dependent correction as $e^{-\frac{ap_{0}}{\hbar}}$.
We have also showed that the commutators of the $\kappa$-deformed gauge covariant derivative, $([\tilde{D}_{0},\tilde{D}_{i}]\phi)^{a}$ and $([\tilde{D}_{i},\tilde{D}_{j}]\phi)^{a}$ match with the Yang-Mills field strength tensors $\hat{F}_{0i}$ and $\hat{F}_{ij}$, respectively, by explicit calculations. We have constructed the Lagrangian describing Yang-Mills theory in $\kappa$-space-time using the field strength tensor and two of the Yang-Mills field equation are derived by finding the corresponding Euler-Lagrange equation. We analysed the symmetry of the system and showed that the Lagrangian is invariant under SU(N) transformation. We have verified that in the commutative limit, i.e., $a\rightarrow 0$, we get the commutative results, as expected.
8 Acknowledgement
BR thanks DST-INSPIRE for support through the INSPIRE fellowship (IF220179).
Data availability statement
No new data were created or analyzed in this study.
Conflict of interest
The authors declare no Conflict of interest.